OpenStudy (anonymous):

integral of 1/sqrt{2}

OpenStudy (dumbcow):

\[\int\limits_{}^{}\frac{1}{\sqrt{2}} = \int\limits_{}^{}2^{-1/2} = 2*2^{1/2} =2\sqrt{2}\]